# SOUL: Unlocking the Power of Second-Order Optimization for LLM Unlearning

### v3

@@ Abstract
Large Language Models (LLMs) have highlighted the necessity of effective unlearning mechanisms to comply with data regulations and ethical AI practices. LLM unlearning aims at removing undesired data influences and associated model capabilities without compromising utility out of the scope of unlearning. While interest in studying LLM unlearning is growing, the impact of the optimizer choice for LLM unlearning remains under-explored. In this work, we shed light on the significance of optimizer selection in LLM unlearning for the first time, establishing a clear connection between second-order optimization and influence unlearning (a classical approach using influence functions to update the model for data influence removal). This insight propels us to develop a second-order unlearning framework, termed SOUL, built upon the second-order clipped stochastic optimization (Sophia)-based LLM training method. SOUL extends the static, one-shot model update using influence unlearning to a dynamic, iterative unlearning process. Our extensive experiments show that SOUL consistently outperforms conventional first-order methods across various unlearning tasks, models, and metrics, suggesting the promise of second-order optimization in providing a scalable and easily implementable solution for LLM unlearning. Codes are available at https://github.com/OPTML-Group/SOUL.
@@ 1 Introduction
LLMs have emerged as transformative technology, greatly enhancing natural language processing capabilities from text generation to simulating human-like interactions Bubeck et al. (2023); Touvron et al. (2023). While offering substantial benefits, LLMs also present significant challenges, such as the risk of misuse in generating private, toxic, or illegal content Nasr et al. (2023); Wen et al. (2023); Karamolegkou et al. (2023); Sun et al. (2024), perpetuation of biases Motoki et al. (2023); Kotek et al. (2023), and the potential for aiding in developing cyberattacks or bioweapons Barrett et al. (2023); Li et al. (2024b).
@@
To address the aforementioned risks, the problem of LLM unlearning arises, aimed at eliminating specific undesirable data influences and their corresponding model generation capabilities while ensuring that model utility is not compromised out of the unlearning scope Jang et al. (2022); Wang et al. (2023); Chen and Yang (2023); Yao et al. (2023); Eldan and Russinovich (2023); Yao et al. (2024); Liu et al. (2024b); Li et al. (2024b); Zhang et al. (2024). While the concept is appealing, the development of effective unlearning algorithms remains challenging. A straightforward approach involves retraining the model from scratch after removing the undesired training data, driven by data privacy concerns Nguyen et al. (2022); Thudi et al. (2022). However, this method is impractical due to the extremely high cost associated with retraining LLMs from scratch. Therefore, model fine-tuning under a predefined unlearning objective has become the primary approach to solve most LLM unlearning problems Jang et al. (2022); Yao et al. (2023); Eldan and Russinovich (2023); Maini et al. (2024). Unfortunately, there is a lack of effective fine-tuning techniques for LLM unlearning. For example, classical gradient ascent-based fine-tuning techniques are susceptible to over-forgetting, which can hamper the original model utility Yao et al. (2023); Maini et al. (2024). Conversely, less aggressive fine-tuning techniques, such as fine-tuning solely on the retain set (i.e., the data set irrelevant to the forgetting data points) Yao et al. (2023), could result in under-forgetting, failing to completely erase the influence of forgotten data. As a result, it is hard to strike the optimal balance between unlearning effectiveness and model utility preservation.
@@
Several recent efforts have been made to develop improved model fine-tuning techniques for LLM unlearning. For example, studies have delved into designing fine-tuning loss functions tailored for LLM unlearning Yao et al. (2023); Eldan and Russinovich (2023); Maini et al. (2024). A currently popular choice is the regularized optimization objective that integrates unlearning efficacy loss with model utility loss, as seen in approaches such as the gradient difference (GradDiff) Yao et al. (2023); Maini et al. (2024) and preference optimization (PO) Eldan and Russinovich (2023); Maini et al. (2024). Additionally, other LLM unlearning techniques incorporate the model’s prior into fine-tuning. For instance, fine-tuning is selectively applied to a subset of model units deemed essential for the unlearning task Yu et al. (2023); Wu et al. (2023b). This approach has led to the emergence of the localization-informed LLM unlearning. Furthermore, input prompt-based strategies have been employed, enabling unlearning through model queries and/or adjusting only a small fraction of learnable parameters Madaan et al. (2022); Zheng et al. (2023); Pawelczyk et al. (2023).
@@
Despite the recent progress of LLM unlearning, the majority of existing fine-tuning-based approaches have relied on first-order (FO) optimization to conduct unlearning. To our knowledge, there have been no prior studies that specifically investigate LLM unlearning from the perspective of optimizer design. In this work, we unveil the power of second-order (SO) optimizer in LLM unlearning and demonstrate its superiority over FO optimizer under the same fine-tuning objective. We will show that SO optimization not only offers a viable approach for enhancing unlearning efficacy but also stays effective in preserving model utility. Such an optimizer-induced advantage holds consistently across various LLM unlearning objectives and formulations, providing a generic improvement. We summarize our contributions below.
@@
Figure 1: Performance highlight using SO optimization in the TOFU dataset Maini et al. (2024) for fictitious unlearning. (Left) Examples of text outputs from LLMs post unlearning using various approaches, including FO GradDiff and PO Yao et al. (2023); Maini et al. (2024); Eldan and Russinovich (2023), as well as their SO counterparts. Failed unlearning is indicated by undesired answers marked in red, while successful unlearning is highlighted in green for desired answers. (Right) Quantitative evaluation comparing SO unlearning with FO unlearning using the metrics forget quality and model utility, as detailed in Sec. 5.
@@
$\bullet$ We study the impact of optimizer choice in LLM unlearning, explicitly linking SO optimization and influence unlearning, which utilizes the influence function approach for unlearning.
@@
$\bullet$ We propose SOUL, a novel SO LLM unlearning framework, which is built upon and extended from Sophia (second-order clipped stochastic optimization) Liu et al. (2023a). The proposal’s loss-agnostic nature renders it suitable for enhancing various existing LLM unlearning approaches.
@@
$\bullet$ We conduct extensive experiments across various LLM unlearning tasks, models, and evaluation metrics, consistently showing the effectiveness of SOUL in improving LLM unlearning, as highlighted in Fig. 1.
@@ 2 Related Work
The concept of machine unlearning has emerged from data protection regulations, such as the ‘right to be forgotten’ Rosen (2011), which were initially not specifically targeted at LLMs Cao and Yang (2015); Hoofnagle et al. (2019); Bourtoule et al. (2021); Nguyen et al. (2022).
As the field has progressed, the applications of machine unlearning have rapidly expanded into diverse areas such as image classification Ginart et al. (2019); Golatkar et al. (2020); Kurmanji et al. (2023); Jia et al. (2023), text-to-image and image-to-image generation Gandikota et al. (2023); Zhang et al. (2023b); Kumari et al. (2023); Fan et al. (2024b); Li et al. (2024a), federated learning Liu et al. (2020); Wang et al. (2022); Liu et al. (2023b), and graph neural networks Chen et al. (2022); Wu et al. (2023a).
@@
The exploration of machine unlearning in the context of LLMs has garnered increasing interest Jang et al. (2022); Wang et al. (2023); Chen and Yang (2023); Yao et al. (2023); Eldan and Russinovich (2023); Yao et al. (2024); Liu et al. (2024b); Li et al. (2024b); Zhang et al. (2024). Seminal works by Liu et al. (2024a) and Zhang et al. (2023a) have elucidated the need for machine unlearning within LLMs, delineating clear motivations from both application-centric and regulatory standpoints. Some research efforts Jang et al. (2022); Yao et al. (2023); Chen and Yang (2023); Maini et al. (2024); Zhang et al. (2024) have concentrated on employing gradient ascent to facilitate forgetting in targeted datasets. Other studies such as those by Maini et al. (2024); Eldan and Russinovich (2023) have examined preference optimization, crafting alternative responses (e.g., reject) to realize unlearning. In addition, some unlearning methods have explored and exploited the data-model interactions that could affect LLM unlearning Meng et al. (2022); Yu et al. (2023); Wu et al. (2023b), such as weight localization-informed unlearning Yu et al. (2023), and altering the hidden representations of LLMs to achieve unlearning Li et al. (2024b). Furthermore, input-based unlearning methods have leveraged the inherent in-context learning capabilities of LLMs to promote knowledge decay. For instance, Thaker et al. (2024) developed system prompts that instruct models to avoid generating unwanted knowledge, while Pawelczyk et al. (2023) applied in-context learning strategies to address unlearning. Last but not least, some recent benchmarks have been developed for the evaluation of LLM unlearning, such as TOFU for fictitious unlearning Maini et al. (2024) and WMDP for unlearning hazardous knowledge in LLMs Li et al. (2024b). Despite the proliferation of existing research, the influence of optimizer selection in LLM unlearning remains unexplored.
3 A Primer on LLM Unlearning
In this section, we outline the LLM unlearning problem setup and discuss common unlearning formulations. These will serve as the starting points and baselines for our proposal.
@@ 3 A Primer on LLM Unlearning
The concept of LLM unlearning Eldan and Russinovich (2023); Yao et al. (2023); Maini et al. (2024); Liu et al. (2024a), aims to mitigate the influence of undesired data, such as sensitive or copyrighted information, and/or restrict the model’s capabilities, such as avoid harmful content generation. This should also be achieved while preserving the LLM’s utility for unrelated tasks and avoid full retraining for computation efficiency.
@@
While problem (2) may appear as a straightforward optimization task initially, complexities arise in determining the effective forget loss $\ell_{\mathrm{f}}$ and achieving the optimal balance between unlearning and utility. These questions remain challenging in the literature. We present two representative LLM unlearning approaches and illustrate how they relate to the specifics of problem (2).
(a) Gradient Difference (GradDiff) Yao et al. (2023); Maini et al. (2024). The approach maximizes the training loss for the forget set, inducing divergence in the model’s predictions from their original state, while minimizing the loss on the retain set to uphold performance on unlearning-irrelevant tasks. Let $\ell(y|x;\bm{\theta})$ denote the prediction loss of using the model $\bm{\theta}$ given the input $x$ against the undesired response $y$. Then, the forget loss $\ell_{\mathrm{f}}$ can be specified by utilizing the negative training loss over the forget set $\mathcal{D}_{\mathrm{f}}$, while the retain loss remains the same as the training loss. This specifies (2) as
@@
Evaluation of LLM unlearning.
While evaluation methods and metrics for LLM unlearning may vary depending on the specific LLM tasks, their main goal is to determine whether unlearning approaches can effectively remove the influence of the unlearning target in LLM generation without compromising unrelated functionalities Liu et al. (2024a). Please refer to Sec. 5.1 for further details on the evaluation metrics utilized.
4 Harnessing Second-Order Optimization to Enhance LLM Unlearning: Why and How
In this section, we shed light on a missing factor of LLM unlearning: the choice of optimizer, which has been overlooked in the literature yet crucial for the effectiveness of unlearning. We advocate for the adoption of a SO (second-order) optimizer for LLM unlearning, drawing inspiration from a classic influence unlearning approach commonly utilized in vision tasks. This stands in contrast to the conventional default FO (first-order) gradient-based optimizer for solving LLM unlearning problems, such as GradDiff and PO introduced in Sec. 3. Based on our insight, we propose SOUL, a second-order unlearning framework for LLMs.
Revisiting influence unlearning and gaining insights.
@@ 4 Harnessing Second-Order Optimization to Enhance LLM Unlearning: Why and How
$\bm{\theta}_{\mathrm{o}}=\bm{\theta}(\mathbf{1}),~{}~{}\bm{\theta}_{\mathrm{MU}}=\bm{\theta}(\mathbf{w}_{\mathrm{MU}}),$ (8)
@@
The derivation of (9) is highly non-trivial as the retrained model $\bm{\theta}(\mathbf{w}_{\mathrm{MU}})$ cannot be directly obtained and is implicitly defined through the optimization problem $\min_{\bm{\theta}}\ell(\bm{\theta},\mathbf{w}_{\mathrm{MU}})$ in (7). To proceed, the influence function approach Koh and Liang (2017); Grosse et al. (2023); Jia et al. (2023) simplifies (9) by applying a first-order Taylor expansion to $\bm{\theta}(\mathbf{w}_{\mathrm{MU}})$ at $\mathbf{w}=\mathbf{1}$:
@@
$\approx$ $\frac{d\bm{\theta}(\mathbf{w})}{d\mathbf{w}}\left|\right._{\mathbf{w}=\mathbf{1}}(\mathbf{w}_{\mathrm{MU}}-\mathbf{1}),$ (10)
where $\frac{d\bm{\theta}(\mathbf{w})}{d\mathbf{w}}$ denotes the full derivative of $\bm{\theta}(\mathbf{w})$ with respect to (w.r.t.) $\mathbf{w}$, and is known as implicit gradient Gould et al. (2016); Zhang et al. (2023d). Utilizing the implicit function theorem Krantz and Parks (2002), the closed form of the influence unlearning formula (10) can be derived as below (Jia et al., 2023, Proposition 1):
@@
where $\ell(\bm{\theta},\mathbf{w})$ represents the $\mathbf{w}$-weighted training loss (7), $\mathbf{H}^{-1}$ stands for the inverse of the second-order derivative (i.e., Hessian matrix) $\nabla_{\bm{\theta},\bm{\theta}}\ell(\bm{\theta},\mathbf{1}/N)$ evaluated at $\bm{\theta}_{\mathrm{o}}$, $\nabla_{\bm{\theta}}\ell$ denotes the gradient of $\ell$, and $\mathbf{1}-\mathbf{w}_{\mathrm{MU}}$ captures the data weight change from pre-training to unlearning encoded in $\mathbf{w}_{\mathrm{MU}}$. To compute (11), one must determine the inverse-Hessian gradient product. However, exact computation is often computationally prohibitive. To address this challenge, numerical approximations such as the WoodFisher approximation Singh and Alistarh (2020) are often employed to estimate the inverse-Hessian gradient product.
@@
If we can transition from the static, one-shot nature of influence unlearning to a dynamic, iterative optimization process, we anticipate that the diminished accuracy resulting from the approximations used in influence unlearning (11) will be mitigated through the iterative engagement of the learning process. However, we still face the computational challenge posed by the Hessian inversion in (12). Therefore, we need to select a practically feasible SO optimization method for LLM unlearning.
Sophia (Second-order Clipped Stochastic Optimization) Liu et al. (2023a) is well-suited since it utilizes a simple diagonal matrix estimate of the Hessian and has shown its effectiveness in LLM pre-training. Sophia modifies the vanilla Newton’s method to
@@
where $\mathbf{m}_{t}\leftarrow\beta_{1}\mathbf{m}_{t-1}+(1-\beta_{1})\mathbf{g}_{t}$ is the exponential moving average (EMA) of the FO gradient with parameter $\beta_{1}>0$, $\mathbf{h}_{t}$ denotes the EMA of the Hessian diagonal estimates obtained from the diagonal of the Gauss-Newton matrix Liu et al. (2023a), and the clipping operation $\mathrm{clip}(\bm{\theta},a)$ limits the magnitude of each element in vector $\bm{\theta}$ to a maximum of $a$, thereby preventing excessively large updates that could destabilize the optimization process. In (13), both the clipping operation $\mathrm{clip}(\cdot,\cdot)$ and the division operation $\cdot$/$\cdot$ / ⋅ are all performed element-wise, and $\gamma>0$ and $\epsilon>0$ are additional parameters in the clipping operation. In (13), if the clipping operation is absent with $\gamma=1$ and $\epsilon\to 0$, then the Sophia update (13) simplifies to the Newton update (12) utilizing the diagonal Hessian estimate for $\mathbf{H}$. In what follows, we will adopt Sophia as the default SO optimizer.
We next link influence unlearning (11) with the SO optimizer and propose the SO unlearning approach. Recall from (11) and (7) that the change in data weights $(\mathbf{1}-\mathbf{w}_{\mathrm{MU}})$ encodes the influence of the forget set $\mathcal{D}_{\mathrm{f}}$ in model training. Therefore, we can interpret the term $\mathbf{H}^{-1}\nabla_{\bm{\theta}}\ell(\bm{\theta}_{0},\mathbf{1}-\mathbf{w}_{\mathrm{MU}})$ in (11) as a SO optimization-based ascent step over the forget set. This is different from the original Sophia update (13), which executes the descent Newton step. Let us take GradDiff (4) as an example. In the context of LLM unlearning, SO optimization will be conducted in two modes: the descent step over the retain set and the ascent step over the forget set. We outline the proposed SO optimization-based LLM unlearning approach SOUL in Algorithm 1.
Algorithm 1 SOUL to solve problem (4)
2:for $t=1$ to $T$ do
7:     Based on $\mathbf{m}_{t}$ and $\mathbf{h}_{t}$, update $\bm{\theta}$ based on (13):
$\bm{\theta}_{t}=\left\{\begin{array}[]{r}\bm{\theta}_{t-1}+\eta_{t}\mathrm{clip}(\mathbf{m}_{t}/\mathrm{max}\left\{{\gamma\mathbf{h}_{t}},\epsilon\right\},1)\\ \text{(ascent mode for forget data)}\\ \bm{\theta}_{t-1}-\eta_{t}\mathrm{clip}(\mathbf{m}_{t}/\mathrm{max}\left\{{\gamma\mathbf{h}_{t}},\epsilon\right\},1)\\ \text{(descent mode for retain data)}\end{array}\right.$ (18)
8:end for
When considering PO-type problems like (6), step 7 of Algorithm 1, as depicted in (18), can only operate in the descent mode. This is because the preference (i.e., the unlearning response $y_{\mathrm{f}}$) has already been defined, and the corresponding forget loss is minimized rather than maximized in (4). In this scenario, SOUL enables the optimization of both forget loss and retain loss through descent mode unification.

### v4

@@ Abstract
Large Language Models (LLMs) have highlighted the necessity of effective unlearning mechanisms to comply with data regulations and ethical AI practices. LLM unlearning aims at removing undesired data influences and associated model capabilities without compromising utility beyond the scope of unlearning. While interest in studying LLM unlearning is growing, the impact of the optimizer choice for LLM unlearning remains unexplored. In this work, we shed light on the significance of optimizer selection in LLM unlearning for the first time, establishing a clear connection between second-order optimization and influence unlearning (a classical approach using influence functions to update the model for data influence removal). This insight propels us to develop a second-order optimization-based LLM unlearning framework, termed Second-Order UnLearning (SOUL), which extends the static, one-shot model update using influence unlearning to a dynamic, iterative unlearning process. Our extensive experiments show that SOUL consistently outperforms conventional first-order methods across various unlearning tasks, models, and metrics, indicating that second-order optimization offers an effective and broadly applicable solution for LLM unlearning. Codes are available at https://github.com/OPTML-Group/SOUL.
@@ 1 Introduction
LLMs have emerged as transformative technology, greatly enhancing natural language processing capabilities from text generation to simulating human-like interactions Touvron et al. (2023). While offering substantial benefits, LLMs also present challenges, such as the risk of misuse in generating private, toxic, or illegal content Nasr et al. (2023); Wen et al. (2023); Karamolegkou et al. (2023); Sun et al. (2024), perpetuation of biases Motoki et al. (2023); Kotek et al. (2023), and the potential for aiding in developing cyberattacks or bioweapons Barrett et al. (2023); Li et al. (2024b).
@@
To address the aforementioned risks, the problem of LLM unlearning arises, aimed at eliminating specific undesirable data influences and their corresponding model generation capabilities while ensuring that model utility is not compromised out of the unlearning scope Liu et al. (2024a); Jang et al. (2022); Wang et al. (2023); Chen and Yang (2023); Yao et al. (2023); Eldan and Russinovich (2023); Yao et al. (2024); Liu et al. (2024b); Li et al. (2024b); Zhang et al. (2024). While the concept is appealing, the development of effective unlearning algorithms remains challenging. A straightforward approach involves retraining the model from scratch after removing the undesired training data, driven by data privacy concerns Nguyen et al. (2022); Thudi et al. (2022). However, this method is impractical due to the extremely high cost associated with retraining LLMs from scratch. Therefore, model fine-tuning under a predefined unlearning objective has become the primary approach to solve most LLM unlearning problems Jang et al. (2022); Yao et al. (2023); Eldan and Russinovich (2023); Maini et al. (2024). Unfortunately, there is a lack of effective fine-tuning techniques for LLM unlearning. For example, classical gradient ascent-based fine-tuning techniques are susceptible to over-forgetting, which can hamper the original model utility Yao et al. (2023); Maini et al. (2024); Zhang et al. (2024). Conversely, less aggressive fine-tuning techniques, such as fine-tuning solely on the retain set (i.e., the data set irrelevant to the forgetting data points) Yao et al. (2023), could result in under-forgetting, failing to completely erase the influence of forgotten data. As a result, it is hard to strike the optimal balance between unlearning effectiveness and model utility preservation.
@@
Several recent efforts have been made to develop improved model fine-tuning techniques for LLM unlearning. For example, studies have delved into designing fine-tuning loss functions tailored for LLM unlearning Yao et al. (2023); Eldan and Russinovich (2023); Zhang et al. (2024). A currently popular choice is the regularized optimization objective that integrates unlearning efficacy loss with model utility loss, as seen in approaches such as the gradient difference (GradDiff) Liu et al. (2022); Yao et al. (2023); Maini et al. (2024), preference optimization (PO) Eldan and Russinovich (2023); Maini et al. (2024) and negative preference optimization (NPO) Zhang et al. (2024). Additionally, other LLM unlearning techniques incorporate the model’s prior into fine-tuning. For instance, fine-tuning is selectively applied to a subset of model units deemed essential for the unlearning task Yu et al. (2023); Wu et al. (2023). This approach has led to the emergence of localization-informed LLM unlearning Liu et al. (2024a). Furthermore, input prompt strategies have been employed, enabling unlearning through model queries and/or adjusting only a small fraction of parameters Madaan et al. (2022); Zheng et al. (2023); Pawelczyk et al. (2023).
@@
Despite the recent progress of LLM unlearning, the majority of existing fine-tuning-based approaches have relied on first-order (FO) optimization to conduct unlearning. To our knowledge, there have been no prior studies that specifically investigate LLM unlearning from the perspective of optimizer design. In this work, we unveil the power of second-order (SO) optimizer in LLM unlearning and demonstrate its superiority over FO optimizer in various fine-tuning scenarios. We term the second-order optimization-based unlearning framework as SOUL (second-order unlearning). We will show that SOUL not only offers a viable approach for enhancing unlearning efficacy but also stays effective in preserving model utility. Such an optimizer-induced advantage holds consistently across various LLM unlearning objectives and formulations, providing a generic improvement. We summarize our contributions below.
@@
Figure 1: Performance highlight using SO optimization (SOUL) in the TOFU dataset Maini et al. (2024) for fictitious unlearning. (Left) Examples of text outputs from LLMs post unlearning using various approaches, including FO GradDiff (gradient difference) Liu et al. (2022); Maini et al. (2024) and PO (preference optimization) Maini et al. (2024); Eldan and Russinovich (2023), as well as their SO counterparts. Failed unlearning is indicated by undesired answers marked in red, while successful unlearning is highlighted in green for desired answers. (Right) Quantitative evaluation comparing SO unlearning with FO unlearning using the metrics forget quality and model utility, as detailed in Sec. 5.
@@
$\bullet$ We study the impact of optimizer choice in LLM unlearning, explicitly linking SO optimization and iterative influence unlearning.
@@
$\bullet$ We propose SOUL, built upon and extended from Sophia (second-order clipped stochastic optimization) Liu et al. (2023a). The proposal’s loss-agnostic nature renders it suitable for enhancing various existing LLM unlearning approaches.
@@
$\bullet$ We conduct thorough experiments across various LLM unlearning tasks, models, and evaluation metrics, consistently showing the effectiveness of SOUL in improving LLM unlearning, as exemplified in Fig. 1.
@@ 2 Related Work
The concept of machine unlearning has emerged from data protection regulations, such as the ‘right to be forgotten’ Rosen (2011), which were initially not specifically targeted at LLMs Cao and Yang (2015); Hoofnagle et al. (2019); Bourtoule et al. (2021); Nguyen et al. (2022). As the field has progressed, the applications of machine unlearning have rapidly expanded into diverse areas such as image classification Ginart et al. (2019); Golatkar et al. (2020); Kurmanji et al. (2023); Jia et al. (2023), text-to-image and image-to-image generation Gandikota et al. (2023); Zhang et al. (2023b); Kumari et al. (2023); Fan et al. (2024b); Li et al. (2024a), and federated learning Wang et al. (2022); Liu et al. (2023b).
@@
The exploration of machine unlearning in the context of LLMs has garnered increasing interest Jang et al. (2022); Wang et al. (2023); Chen and Yang (2023); Yao et al. (2023); Eldan and Russinovich (2023); Yao et al. (2024); Liu et al. (2024b); Li et al. (2024b); Zhang et al. (2024). Seminal works by Liu et al. (2024a) and Zhang et al. (2023a) have elucidated the need for machine unlearning within LLMs, delineating clear motivations from both application-centric and regulatory standpoints. Some research efforts Jang et al. (2022); Yao et al. (2023); Chen and Yang (2023); Maini et al. (2024); Zhang et al. (2024) have concentrated on employing gradient ascent to facilitate forgetting in targeted datasets. Other studies such as those by Maini et al. (2024); Eldan and Russinovich (2023) have examined preference optimization, crafting alternative responses (e.g., reject) to realize unlearning. In addition, some unlearning methods have explored and exploited the data-model interactions that could affect LLM unlearning Meng et al. (2022); Yu et al. (2023); Wu et al. (2023), such as weight localization-informed unlearning Yu et al. (2023), and altering the hidden representations of LLMs to achieve unlearning Li et al. (2024b). Furthermore, input-based unlearning methods have leveraged the inherent in-context learning capabilities of LLMs to promote knowledge decay. For instance, Thaker et al. (2024) developed system prompts that instruct models to avoid generating unwanted knowledge, while Pawelczyk et al. (2023) applied in-context learning strategies to address unlearning. Last but not least, some recent benchmarks have been developed for the evaluation of LLM unlearning, such as TOFU for fictitious unlearning Maini et al. (2024) and WMDP for unlearning hazardous knowledge in LLMs Li et al. (2024b). Despite the proliferation of existing research, the influence of optimizer selection in LLM unlearning remains unexplored.
3 Primer on LLM Unlearning
@@
LLM unlearning aims to mitigate the influence of undesired data, such as sensitive or copyrighted information, and/or restrict the model’s capabilities to avoid the associated content generation. This process also requires preserving the LLM’s utility for unrelated tasks and avoiding full retraining to maintain computational efficiency.
@@
While problem (2) may appear as a straightforward optimization task initially, complexities arise in determining the effective forget loss $\ell_{\mathrm{f}}$ and achieving the optimal balance between unlearning and utility. These questions remain challenging in the literature. We present three representative LLM unlearning approaches and illustrate how they relate to the specifics of problem (2).
(a) Gradient Difference (GradDiff) Liu et al. (2022); Maini et al. (2024). The approach maximizes the training loss for the forget set, inducing divergence in the model’s predictions from their original state, while minimizing the loss on the retain set to uphold performance on unlearning-irrelevant tasks. Let $\ell(y|x;\bm{\theta})$ denote the prediction loss of using the model $\bm{\theta}$ given the input $x$ against the undesired response $y$. Then, the forget loss $\ell_{\mathrm{f}}$ can be specified by utilizing the negative training loss over the forget set $\mathcal{D}_{\mathrm{f}}$, while the retain loss remains the same as the training loss. This specifies (2) as
@@
(c) Negative Preference Optimization (NPO) Zhang et al. (2024). NPO also treats the unlearning problem as a preference optimization problem. Yet, different from PO that specifies the unlearning response $y_{\mathrm{f}}$, it interprets the forgetting data in $\mathcal{D}_{\mathrm{f}}$ as the negative examples and incorporates them alone in preference optimization Rafailov et al. (2024). This yields a similar problem as GradDiff (4), but replaces the GA loss with the negative examples-based preference optimization loss.
4 Second-Order Optimization to Enhance LLM Unlearning: Why & How
In this section, we shed light on a missing factor of LLM unlearning: the choice of optimizer, which has been overlooked in the literature yet crucial for the effectiveness of unlearning.
Gaining insights from influence unlearning.
@@ 4 Second-Order Optimization to Enhance LLM Unlearning: Why & How
$\bm{\theta}_{\mathrm{o}}=\bm{\theta}(\mathbf{1}),~{}~{}\bm{\theta}(\mathbf{w}_{\mathrm{MU}})=\bm{\theta}_{\mathrm{MU}},$ (8)
@@
The derivation of (9) is highly non-trivial as the retrained model $\bm{\theta}(\mathbf{w}_{\mathrm{MU}})$ cannot be directly obtained and is implicitly defined through the optimization problem $\min_{\bm{\theta}}\ell(\bm{\theta},\mathbf{w}_{\mathrm{MU}})$. To proceed, the influence function approach Koh and Liang (2017); Grosse et al. (2023); Jia et al. (2023) simplifies (9) by applying a first-order Taylor expansion to $\bm{\theta}(\mathbf{w}_{\mathrm{MU}})$ at $\mathbf{w}=\mathbf{1}$:
@@
$\approx$ $\frac{d\bm{\theta}(\mathbf{w})}{d\mathbf{w}}\left|\right._{\mathbf{w}=\mathbf{1}}(\mathbf{w}_{\mathrm{MU}}-\mathbf{1}),\vspace*{-3mm}$ (10)
where $\frac{d\bm{\theta}(\mathbf{w})}{d\mathbf{w}}$ denotes the full derivative of $\bm{\theta}(\mathbf{w})$ with respect to (w.r.t.) $\mathbf{w}$, and is known as implicit gradient Gould et al. (2016); Zhang et al. (2023d). Utilizing the implicit function theorem Krantz and Parks (2002), the closed form of the influence unlearning formula (10) can be given by (Jia et al., 2023, Proposition 1):
@@
where $\ell(\bm{\theta},\mathbf{w})$ represents the $\mathbf{w}$-weighted training loss (7), $\mathbf{H}^{-1}$ stands for the inverse of the second-order derivative (i.e., Hessian matrix) $\nabla_{\bm{\theta},\bm{\theta}}\ell(\bm{\theta},\mathbf{1}/N)$ evaluated at $\bm{\theta}_{\mathrm{o}}$, $\nabla_{\bm{\theta}}\ell$ denotes the gradient of $\ell$, and $\mathbf{1}-\mathbf{w}_{\mathrm{MU}}$ yields $\mathbf{1}-\mathbf{1}_{\mathcal{D}_{\mathrm{r}}}$, which captures the data weight on the forget set $\mathcal{D}_{\mathrm{f}}$. To compute (11), one must determine the inverse-Hessian gradient product. However, exact computation is often computationally prohibitive. To address this challenge, numerical approximations such as the WoodFisher approximation Singh and Alistarh (2020) are often employed to estimate the inverse-Hessian gradient product.
@@
If we can transition from the static, one-shot nature of influence unlearning to a dynamic, iterative optimization process, we anticipate that the diminished accuracy resulting from the approximations used in influence unlearning (11) will be mitigated through the iterative engagement of the learning process. However, we still face the computational challenge posed by the Hessian inversion in (12). Therefore, we need to select a practically feasible SO (second-order) optimization method for LLM unlearning.
Sophia (Second-order Clipped Stochastic Optimization) Liu et al. (2023a), a simple scalable SO optimizer, is well-suited since it utilizes a simple diagonal matrix estimate of the Hessian and has shown its effectiveness in LLM pre-training. Sophia modifies the vanilla Newton’s method to
@@
where $\mathbf{m}_{t}\leftarrow\beta_{1}\mathbf{m}_{t-1}+(1-\beta_{1})\mathbf{g}_{t}$ is the exponential moving average (EMA) of the FO (first-order) gradient with parameter $\beta_{1}>0$, $\mathbf{h}_{t}$ denotes the EMA of the Hessian diagonal estimates obtained from the diagonal of the Gauss-Newton matrix Liu et al. (2023a), and the clipping operation $\mathrm{clip}(\bm{\theta},a)$ limits the magnitude of each element in vector $\bm{\theta}$ to a maximum of $a$, thereby preventing excessively large updates that could destabilize the optimization process. In (13), both the clipping operation $\mathrm{clip}(\cdot,\cdot)$ and the division operation $\cdot$/$\cdot$ / ⋅ are all performed element-wise, and $\gamma>0$ and $\epsilon>0$ are additional parameters in the clipping operation. In (13), if the clipping operation is absent with $\gamma=1$ and $\epsilon\to 0$, then the Sophia update (13) simplifies to the Newton update (12) utilizing the diagonal Hessian estimate for $\mathbf{H}$.
Next, we link influence unlearning (11) with the SO optimizer and propose the SO unlearning approach. Recall from (11) and (7) that the change in data weights $(\mathbf{1}-\mathbf{w}_{\mathrm{MU}})$ encodes the influence of the forget set $\mathcal{D}_{\mathrm{f}}$ in model training. Therefore, we can interpret the term $\mathbf{H}^{-1}\nabla_{\bm{\theta}}\ell(\bm{\theta}_{0},\mathbf{1}-\mathbf{w}_{\mathrm{MU}})$ in (11) as a second-order optimization-based ascent step over the forget set. This contrasts with the original Sophia update (13), which executes the descent using the clipped Newton step. Let us take GradDiff (4) as an example. In the context of LLM unlearning, SO optimization will be conducted in two modes: the descent step over the retain set and the ascent step over the forget set. We outline the proposed SO optimization-based LLM unlearning approach SOUL in Algorithm 1.
When considering PO-type problems like (6), the proposed algorithm can only operate in the descent mode. This is because the preference (i.e., the unlearning response $y_{\mathrm{f}}$) has already been defined, and the corresponding forget loss is minimized rather than maximized in (4). In this scenario, SOUL enables the optimization of both forget loss and retain loss through descent mode unification.